OpenStudy (anonymous):

( 7 + radical 3)( 7 - radical 3)

OpenStudy (anonymous):

(7+sqrt(3))(7-sqrt(3)) 49 -7sqrt(3) + 7sqrt(3) -sqrt(3)sqrt(3) 49 - 3 = 46

OpenStudy (anonymous):

Just multiply like normal (FOILing): \[(7+\sqrt{3})(7-\sqrt{3})\] \[(49+7\sqrt{3}-7\sqrt{3}-3)\] \[49-3\] \[46\] Notice how the radicals actually cancel each other.

OpenStudy (anonymous):

this is a differnce of sqaures, which means that that there will be no middle terms 1) when you foil this out the answer will be 49-3 and your answer will be 46 hope that helps :)